# Digital currency hardware wallets and the essence of money

### v3

@@ Abstract
Many proposals for the design and implementation of digital wallets assume that the purpose of the wallet is to enable offline payments via custodial accounts, ignoring the real problems faced by individuals and businesses that engage in retail payments, such as the anticompetitive behaviour of payment platforms and the decline of cash. More importantly, the proposals ignore the raison d’être of digital currency as a kind of digital money that can be held independently of custodians. Finally, the proposals demonstrate a profound lack of imagination about the nature of digital money and the devices that could be used to hold, manage, and exchange it. From these presumptions flows a set of architectural requirements that stifle the promise of digital currency to deliver novel and efficient ways to exchange value in the digital economy. In this article, we critically assess the essential problems that digital currency solutions are being proposed to solve, particularly with respect to the future of payments and the future of cash. We assess the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs. We conclude that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
@@ 1 Introduction
The proliferation of cashless payments has given rise to a growing preponderance of merchants who refuse to accept cash, as well as the suggestion that private-sector payment systems could subvert the authority of central banks to implement monetary policy. This development has, in turn, focussed the attention of central banks and other financial institutions around the world on digital currency, particularly the idea that central banks could issue central bank digital currency (CBDC) for retail use. In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
@@
Many proponents of retail CBDC have argued that it must be “cash-like” to deliver value to end-users. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the assets that they use to engage with the economy. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable dual-offline transactions, namely those without connectivity of any kind between either of the counterparties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. We suggest that this argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that the main promise of digital currency concerns accessing custodial accounts whilst offline. But this assumption ignores decades of history of digital currency, wherein the most groundbreaking systems, including not only Bitcoin but also DigiCash two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step.
@@
Dual-offline transactions impose additional security requirements to prevent double-spending attacks, in addition to the new security requirements applicable to digital bearer instruments in general. Some prominent experts have proposed “trusted computing” [1, 2] as a solution to the security problem, reasoning that if the security of bearer instruments is a cause for concern, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must meet a certain standard. After all, many users are accustomed to carrying certified hardware, and the prevalence of certified hardware among common user devices might be seen as a justification for their use in the context of digital currency. For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [3]. Mobile phones often contain secure enclaves to restrict what their users can do with their devices, for example to prevent users from copying multimedia files.
@@
We also consider the conflation of technology, which can be freely and directly deployed by an individual or business to perform a function, with services, which are functions performed by third parties. In a world of smartphone applications for which most of the computation is nearly always performed by remote service providers, it is easy to ignore this distinction. However, since the ability to possess and control money directly is essential to its function, we must be careful to distinguish between technology that is under the control of the end-user and technology that is under the control of a third party.
@@
Of course, the fact that certain methods are commonly used in some contexts does not justify their use in others, and it certainly does not justify mandating the use of such methods in the context of future infrastructure that intrinsically must be available for general use by the public. It is all too easy for those in the payments industry to assume that everyone has or should have a smartphone and a bank account. Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim.
@@
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [42]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [5]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that trusted computing is an indispensable enabler of digital currency solutions, that digital currency must be held in accounts, that wallets must be linked to the identity of a user, and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [5].
@@ 2 The sophistry of dual-offline payments
Recently, several prominent organisations, including the International Monetary Fund [6] as well as central banks such as the Bank of England [7] and the Bank of Canada [8], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
@@ 2.1 Kinds of payments
  • E-commerce transactions. Transactions that involve consumers using personal computers or mobile devices to interact with websites require the internet by definition. They represent a significant and growing share of the total value consumers spend in the economy, and naturally, they require consumers to be online.
@@
  • Bank transfers. Bank transfers are mechanisms for transferring money from a sender’s account with a sending bank to a recipient’s account with a receiving bank. As account-holders, both the sender and the recipient must be identified. Bank transfers can be arranged online, or an account-holder can arrange a bank transfer via a conversation with a bank teller or a computer controlled by the bank. The sending bank debits the sender’s account, and the receiving bank credits the recipient’s account. Settlement between the banks is managed via an interbank clearing and settlement network. In most cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and bank transfers are usually not considered appropriate at the point of sale or for transactions that require acceptance in real-time.
@@
  • Electronic card payments at the point of sale (POS). Electronic card payments provide a mechanism by which merchants can accept in-person payments from consumers in real-time. The consumer’s bank (the issuer) and the merchant’s bank (the acquirer) both participate in a payment network. The issuer provides a card to the consumer, and the acquirer offers to connect a merchant’s POS device to the payment network. The card serves as a form of identification: Its purpose is to specify a particular bank account and to identify its user as the beneficiary of that account. To protect against unauthorised copying of the identification credential, cards are usually designed with features that provide tamper-resistance. The client is not assumed to have an Internet connection, and the POS device provides a way for the card to authenticate itself to the payment network. If the authentication is successful, then the payment network effectively underwrites the transaction, implicitly providing credit to the issuer and prompting the acquirer to accept the transaction immediately. The payment network informs the merchant that the transaction is approved, and the merchant can then accept the payment in real-time. Merchants pay interchange fees to support this mechanism.
@@
  • Closed-loop payment systems at the point of sale. With closed-loop payment networks, both consumers and merchants are primary customers of a platform operator, which maintains accounts with balances for all of its customers. When a consumer pays a merchant in-person, the platform operator debits the consumer’s account and credits the merchant’s account. For a transaction to succeed, either the consumer or the merchant must have an active connection to the platform operator. The consumer must authenticate successfully to the platform operator, although this can be done via the merchant’s device, as it is with card payment networks. Crucially, banks are not generally involved in transactions; the platform operator manages the process.
@@
Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a dramatic decline in use throughout the world over the past decade.
@@
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [9]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which consumers purchase e-money tokens offline and consummate the transactions offline by selectively revealing proofs of purchase [10]. And so, it is fair to ask: Might the idea of dual-offline payments constitute a solution in search of a problem, an ephithet that a report by the UK House of Lords prominently ascribed to CBDC in general [11]?
2.2 Digital currency use cases

### v4

@@ Abstract
Many proposals for the design and implementation of digital wallets assume that the purpose of the wallet is to enable offline payments via custodial accounts, ignoring the real problems faced by individuals and businesses that engage in retail payments, such as the anticompetitive behaviour of payment platforms and the decline of cash. More importantly, the proposals ignore the raison d’être of digital currency as a kind of digital money that can be held independently of custodians. Finally, the proposals demonstrate a profound lack of imagination about the nature of digital money and the devices that could be used to hold, manage, and exchange it. From these presumptions flows a set of architectural requirements that stifle the promise of digital currency to deliver novel and efficient ways to exchange value in the digital economy. In this article, we critically assess the essential problems that digital currency solutions are being proposed to solve, particularly with respect to the future of payments and the future of cash. We evaluate the validity of common justifications for account-based payments and certified hardware in the context of alternative designs, limitations, and trade-offs. We conclude, referencing some specific designs, that the interests of consumers would be better served by design approaches to digital currency that anticipate that digital assets would be held outside accounts, stored offline, but transacted online, without requiring the use of trusted hardware.
@@ 1 Introduction
The proliferation of cashless payments has given rise to a growing number of merchants who refuse to accept cash, a well-documented and legal practice in some jurisdictions around the world [De Nederlandsche Bank 2023, Grabar 2018] that has correspondingly given rise to government reports and consultations [House of Commons 2025, Australia 2024]. Combined with the emergence of novel private-sector payment systems, the decline of cash subverts the authority of central banks to implement monetary policy. This development has, in turn, focussed the attention of central banks and other financial institutions around the world on digital currency, particularly the idea that central banks could issue central bank digital currency (CBDC) for retail use. In principle, retail CBDC would be a direct obligation of the central bank, like cash or central bank reserves, but existing in digital form and held by ordinary consumers and non-financial businesses.
@@
Some proponents of retail CBDC, including central banks, have argued that it should have cash-like properties to deliver value to end-users [ECB 2021, Bank of England 2023]. This argument is at least partially right. For example, it might be expected that a cash-like system would afford users privacy by design, resistance to discrimination and profiling, and the ability to possess and control the monetary assets that they use to participate in the economy and satisfy their economic needs. These are certainly desirable properties. However, some authorities have interpreted the argument that a digital payment system must be cash-like to mean that a digital payment system must enable “dual-offline” transactions, namely those without connectivity of any kind between either of the transacting parties and any third parties, in contrast to “single-offline” transactions, wherein a mutually trusted third party is reachable by one of the counterparties. The “dual-offline” argument constitutes a leap of logic that must be examined carefully, starting with its implicit assumptions that any new payment system must be appropriate for all scenarios, that existing solutions such as cash are insufficient to handle payment scenarios in which neither party has access to a network connection, and that there is a specific need for future digital currency infrastructure to enable a kind of digital transaction that is not commonly available to users today.
Importantly, the idea that digital currency transactions must mirror the current system for electronic retail payments demonstrates an assumption that all digital money must take the form of an account offered by some service provider who acts as a custodian, and that the main promise of digital currency concerns accessing such an account for use in dual-offline payments (an “offline account”). Following this logic, a digital currency hardware wallet would be a tool by which an authorised account holder could conduct peer-to-peer transactions between devices in the absence of network connections. To achieve this, the wallet itself must enforce a set of rules, for example, the wallet must track the size of the account balance, and the amount of money that the account holder can spend must be limited to that value.
However, there is another, competing view that maintains that the purpose of digital currency is to provide a digital form of cash (“e-cash”), and this view is supported by past digital currency design proposals over the past four decades. Some of the most groundbreaking systems, including not only Bitcoin [Nakamoto 2009] but also DigiCash [Pitta 1999] two decades earlier, drew attention by introducing a way to hold money outside of custodial accounts. Both of these systems were sometimes described as “cryptocurrency” because they were engineered using cryptographic techniques. Importantly, the tokens that users held were explicitly designed to be cash-like, as can be inferred from their names, and intended to be used for payments and not linked to accounts. In the case of Bitcoin, tokens were created and spent using an immutable ledger (in the form of a blockchain system), and in the case of DigiCash, tokens were created and spent as bearer instruments using a redemption process with a central operator, which is assumed to be responsible and accountable for its issuances, which can be held anonymously by prospective payers. The idea that central banks would never support such an architecture ignores the fact that central banks have always supported the issuance and circulation of tokens outside accounts, in the form of cash. Within this framework, offering cash in digital form is the next logical step. We suggest that the “e-cash” vision of digital currency is largely absent from the discussion of the design of digital currency hardware wallets, and we argue that this neglect is a mistake. Contemporary examples of “e-cash” systems with significant deployments include GNU Taler [GNUnet and Taler] and (de facto) Privacy Pass [Davidson 2018], both of which feature the use of non-custodial wallets to hold tokens for redemption.
@@
In both the “offline account” and “e-cash” use cases, dual-offline transactions impose additional security requirements on the transacting parties, in addition to the security requirements applicable to digital transactions in general. In particular, payers must be prevented from falsely claiming that the money they are using to pay has not previously been spent (a “double-spending attack”). Some prominent experts have proposed “trusted computing” [Trusted Computing Group 2004, Roberts 2021] as a solution to the security problem, reasoning that if the possibility that a payer (or a payer’s device) might misbehave, then perhaps there is benefit to be realised in making an effort to ensure that the devices that can hold them must be able to enforce rules directly, a “hardware root of trust”. After all, many users are accustomed to carrying certified hardware, and the prevalence of certified hardware among common user devices might be seen as a justification for their use in the context of digital currency. For example, smart cards commonly rely upon tamper-resistant chips containing secrets issued by their manufacturers [ETSI 2019]. Mobile phones often contain secure enclaves to restrict what their users can do with their devices, for example to prevent users from copying multimedia files.
E-cash systems are designed to work without tracking the ownership of individual tokens on a ledger. A generic asset “template” containing information such as a public key is first created by a prospective consumer, who then uses a zero-knowledge protocol to request a privacy-preserving certification of the digital asset, such as a blind signature, from the issuer, which then becomes part of the asset. The issuer can now recognise the certified digital asset as valid without recognising its owner or the circumstances of its creation. The asset is stored by the end-user directly and is not recorded on a ledger. Double-spending can be prevented in a few ways, one of which is to assume (a) that the issuer processes all transactions as redemptions, (b) that the issuer maintains a ledger to keep track of the set of redeemed assets that it receives in these transactions, and (c) that the issuer, upon receiving an asset in a transaction, checks the ledger to verify that the asset has not already been redeemed [Chaum 1982, Chaum 2021]. An alternative approach, which we call oblivious management, does not require the issuer to process transactions or maintain a ledger of redeemed tokens. In such a system, the assets carry their own provenance information, linking transaction information they carry to rolled-up aggregation entries, one per time period, on a ledger [Goodell 2023, Friolo 2025]; however, neither the tokens nor the value they represent are represented on the ledger at all. In both cases, issuers might keep accounting records of the tokens that they issue or redeem, but the privacy-preserving certification unlinks the issuance of an asset, where the party requesting a new asset may be identified, from its subsequent payment transaction and redemption, wherein only the payee is known.
@@
We also consider the conflation of technology, which can be freely and directly deployed by an individual or business to perform a function, with services, which are functions performed by third parties. In a world of smartphone applications for which most of the computation is nearly always performed by remote service providers, it is easy to ignore this distinction. However, since the ability to possess and control cash directly is functional to its use, we must be careful to distinguish between technology whose functions are governed by the end-user and technology whose functions are governed by a third party.
@@
Of course, the fact that certified hardware controlled by third parties is commonly used in some contexts does not justify its use in others. It certainly does not justify mandating the use of such methods in the design of future public infrastructure, especially if doing so would compel users to relinquish some of their control to third parties. It is all too easy for those in the payments industry to assume that everyone has or should have a smartphone and a bank account. Mandating the use of certified hardware carries many costs, and the benefits of certified hardware are not as clear as its proponents claim [Giesecke+Devrient 2023, Thales 2020].
In particular, we suggest that much of the conversation in Europe about digital currency wallets has followed from the expectation, perhaps inspired by the digital euro regulation proposed by the European Commission [European Commission 2023], that for an institutionally accepted digital currency in general, and central bank digital currency in particular, the salient feature of cash is that it can be transferred in a peer-to-peer way. We disagree. We argue that the salient feature of cash is that it is a bearer instrument, and because peer-to-peer transactions are not a necessary characteristic of bearer instruments, we argue that it is worth considering protocols that do not rely upon wallets to enforce the rules, but for which the locus of enforcement lies with service providers. This argument has separately been articulated by the GNU Taler project [Grothoff and Dold 2021]. Even if such protocols rely upon at least one party to be online, they can be considered social-welfare enhancing with respect to protocols that rely upon users holding devices that they do not control, particularly given the well-established vulnerabilities in trusted hardware [GNU Taler 2021], and notwithstanding assertions offered by the ECB that some hardware technology, such as embedded secure elements and SIM devices, are more robust against attacks than trusted execution environments [ECB 2025].
@@
The aforementioned considerations raise the question of how end-users will hold digital currency assets, and in particular, what the role of a wallet should be. ISO 22739:2024 defines a wallet as an “application or mechanism used to generate, manage, store or use private keys and public keys or other digital assets” [ISO 22739 2024a]. The simplicity of this definition would seem to understate the potentially complex features required to implement token management and custody, an exploration of which is given by Barresi and Zatti [Barresi 2023]. We address some of the assumptions that Barresi and Zatti make about the design of wallets and suggest different ways of viewing the problem. In particular, we critically address the arguments that dual-offline payments constitute the most important use case for digital currency (see Section 2), that trusted computing is an indispensable enabler of digital currency solutions (see Section 3), that digital currency must be held in accounts (see Section 4), and that a wallet is properly seen as “not a place where you can store cash [but] that helps you keep track of it” [Barresi 2023] (see Section 6). Additionally, we address the argument implicit to some e-cash solutions that issuance and transaction processing must be performed by the same actor (see Section 5). Throughout this article, we shall reference a set of specific designs, including e-cash designs, such as GNU Taler, that are based on bearer instruments, as well as a proposal for how to integrate bearer instruments into a mechanism for using an integrity provider to transfer digital assets obliviously, without maintaining records of the assets, without immediately redeeming them for the value they represent, and without recording any information about the assets or the counterparties involved in transactions [Goodell 2023, Friolo 2025].
@@ 2 The sophistry of dual-offline payments
In addition to the ECB [ECB 2025], several prominent organisations, including the International Monetary Fund [Kiff 2022] as well as other central banks, such as the Bank of England [Bank of England 2022] and the Bank of Canada [2], have released reports suggesting that dual-offline payments are desirable or even essential to the functionality of a future retail CBDC system. Such bold statements are striking, considering the prevailing lack of consensus about what the requirements for digital currency should be in general. In this section, we shall critically explore the argument (or implicit assumption) that the purpose of digital currency is to facilitate access to custodial accounts in a dual-offline payment scenario.
@@ 2.1 Kinds of payments
  • E-commerce transactions. Transactions that involve consumers using personal computers or mobile devices to interact with websites require the internet by definition. They represent a significant and growing share of the total value consumers spend in the economy, and naturally, they require consumers to be online. (An example is a consumer paying for goods or services by inputting card details into a form on a web page.)
@@
  • Bank transfers. Bank transfers are mechanisms for transferring money from a sender’s account with a sending bank to a recipient’s account with a receiving bank. As account-holders, both the sender and the recipient must be identified. Bank transfers can be arranged online, or an account-holder can arrange a bank transfer via a conversation with a bank teller or a computer controlled by the bank. The sending bank debits the sender’s account, and the receiving bank credits the recipient’s account. Settlement between the banks is managed via an interbank clearing and settlement network. In some cases, interbank settlement networks cannot guarantee “immediate” settlement within a few seconds, and, despite the emergence of new open banking architectures, account-to-account “push” payments (in contrast to “pull” payments, such as via card networks) have yet to see global adoption at the point of sale or for transactions that require acceptance in real-time. (An example is a consumer instructing her bank to pay an invoice, either over the phone, in a branch, or with a banking app or website.)
@@
  • Electronic card payments at the point of sale (POS). Electronic card payments provide a mechanism by which merchants can accept in-person payments from consumers in real-time. The consumer’s bank (the issuer) and the merchant’s bank (the card payment acquirer) both participate in a payment network. The issuer provides a card to the consumer, and the acquirer provides a POS device to connect the merchant to the payment network. The card serves as a form of identification: Its purpose is to specify a particular bank account and to identify its user as the beneficiary of that account. To protect against unauthorised copying of the identification credential, cards are usually designed with features that provide tamper-resistance. The client is not assumed to have a real-time connection to an acquiring bank or service provider, and with contemporary ISO 8583 payments [ISO 8583 2023], the POS device provides a way for the card to authenticate itself to the payment network. If the authentication is successful, then the payment not immediately settled, and an inter-bank credit is implicit when the acquirer accepts the transaction. When the transaction is accepted, the acquirer has a credit against the issuer, which owes the transaction amount to the acquirer. In the case of a debit card transaction, the issuing bank customer is debited immediately; in the case of a credit card transaction, the issuing bank customer gets a credit in exchange for incurring a debt against the issuing bank (credit card debt). The payment network informs the merchant that the transaction is approved, and the merchant can then accept the payment in real-time. Merchants pay interchange fees to support this mechanism. (An example is a consumer buying a cup of tea by presenting a card to a merchant POS device.)
@@
  • Closed-loop payment systems at the point of sale. With closed-loop payment networks, both consumers and merchants are primary customers of a platform operator, which maintains accounts with balances for all of its customers. When a consumer pays a merchant in-person, the platform operator debits the consumer’s account and credits the merchant’s account. For a transaction to succeed, either the consumer or the merchant must have an active connection to the platform operator. The consumer must authenticate successfully to the platform operator, although this can be done via the merchant’s device, as it is with card payment networks. Crucially, banks are not generally involved in transactions; the platform operator manages the process. (An example is a consumer using an Alipay or WeChat app to purchase goods or services using e-money credits.)
@@
Payment type payer online? payee online? bearer token? dual-offline?
E-commerce yes yes no no
Bank transfer yes no no no
POS card payment no yes no no
POS closed-loop payment maybe yes no no
Cash no no yes yes
E-token bearer payment maybe maybe yes no
Table 1: Features of different payment types.
@@
Notwithstanding some measure of understandable confusion and inconsistency in the use of the term “offline” accompanying calls for considering offline payments to be an essential feature of digital currency, it can be said that some measure of online technology supports virtually all retail payments today, excepting only payments in cash, which has seen a decline in use throughout the world over the past decade, starting well before the pandemic of 2020–2022. Table 1 summarises the various types of payment, including the proposed “e-token bearer payment”, which is a digital online payment using a bearer instrument that is currently not an option for most consumers. It is a payment type like this that could be considered the next step in the evolution of cash, cash for the digital economy.
Like cash, this new bearer instrument would be possessed and controlled directly by its users, as an assignable obligation of the issuer. The issuer, which could be a central bank, would be responsible for fulfilling this obligation, but it would not necessarily hold collateral corresponding to the total volume of its issuance.
@@
With this in mind, it is reasonable to ask: What problem do proponents of retail CBDC intend to solve, for which dual-offline payments constitute a better solution than online payments involving assets that are held offline? We acknowledge that there are some limited real-world use cases for dual-offline payments, such as payments aboard transportation systems that lack reliable network connectivity throughout the system [1]. However, such examples are rare indeed, owing to the prevalence of network connectivity in locations suitable for digital transactions. Furthermore, we note that payment scenarios similar to this example do not actually require dual-offline payments; their requirements could also be satisfied by time-shifting, in which a consumer purchases e-money tokens online and consummates transactions with a predetermined merchant offline, by selectively sending to the merchant tokens for which the consumer had already transferred control to the merchant [Goodell 2023]. (The consumer and the merchant can use a fair exchange protocol to recover any unspent tokens later.) And so, it is fair to ask the question implicit to the title of a report by the UK House of Lords, “Central bank digital currencies: a solution in search of a problem?” [House of Lords 2021].
2.2 Motivations for digital currency